Assuming you purchased a share of stock for ​$76.00 one year​ ago, sold it today for ​$92.99​, and during the year received a dividend payment of ​$5.05​, calculate the​ following:

a. Income.

b. Capital gain​ (or loss).

c. Total return

​(1) In dollars.

​(2) As a percentage of the initial investment
Answer #1

purchase price , p0 = $76

selling price , p1 = $92.99

dividend , d = $5.05

a)

income = d+ (p1-p0) = 5.05 + (92.99-76) = $22.04

b)

Capital gain = p1-p0 = 92.99-76= 16.99

c)

Total return (in dollars) = d+ (p1-p0) = 5.05 + (92.99-76) = $22.04

Total return (as a percentage) = [d+ (p1-p0)]/p0 = [5.05 + (92.99-76)]/76 = 0.29 or 29%